a store room has an 80% alcohol solution and a 30% alcohol solution. How many milliliters of each should be used to obtain 50 milliliters of a 60% solution?
Answer by stanbon(75887)

Let "x" be the amount of 30% solution.
Then "50-x" is the amount of 80% solution.
Equation:
Amt. of alcohol in 30% + Amt. in 80% = Amt. in 60%
30x +80(50-x) = 60(50)
30x +4000 - 80x =3000
-50x = -1000
x = 20 mm of 30% solution
50-x = 30 mm of 80% solution
Cheers,
Stan H.